# SINDI: an Efficient Index for Approximate Maximum Inner Product Search on Sparse Vectors

### v1

@@ 3.3. Cache Optimization
When the dataset size $\mathcal{D}$ reaches the million scale, the distance array becomes very large. Random access to such a long array leads to frequent cache misses, causing the query performance to be highly memory-bound. Moreover, maintaining a distance array of size $\mathcal{D}$ for every query consumes significant memory resources. To limit the length of the distance array, Sindi employs the Window Switch strategy that restricts the range of vector IDs accessed in a single window. Within each window, the reduced size of the distance array makes the access pattern nearly sequential.
@@ 3.3.1. Window Switch
During index construction, Sindi partitions the dataset $\mathcal{D}$ into contiguous ID segments, referred to as windows. The window size is denoted by $\lambda$ ($0<\lambda\leq\|\mathcal{D}\|$), and the total number of windows $\sigma$ is $\left\lceil\frac{\|\mathcal{D}\|}{\lambda}\right\rceil$. The $w$-th window contains vectors from $\vec{x}_{w\lambda}$ to $\vec{x}_{(w+1)\lambda-1}$, and the window index to which vector $\vec{x}_{i}$ belongs is $\left\lfloor\frac{i}{\lambda}\right\rfloor$. Each inverted list is partitioned in the same way, so every list has $\sigma$ windows. We denote the $w$-th window of the $j$-th inverted list by $I_{j,w}$, with $0\leq w<\sigma$.
At query time, the length of the distance array $A$ is set to the window size $\lambda$, and all windows share the same $A$. During a window search, each vector $\vec{x}_{i}$ is mapped to a unique entry $A[i\bmod\lambda]$. The search procedure in the $w$-th window consists of two steps: (1) Inner product computation. For each scanned list, compute inner products for vectors $\vec{x}_{w\lambda}$ to $\vec{x}_{(w+1)\lambda-1}$. The computation follows the same two-stage process as described in Section 3.2, namely, product computation followed by accumulation. (2) Heap update. After computing the inner product for the current window, the distance array $A$ contains the final distance. Scan $A$ to insert the top candidates (with the largest inner products) into a minimum heap $H$, which maintains the vector IDs and distances of the results to be returned. Note that we need to recover the vector ID from $A$’s index, $A[t]$ corresponds to $\vec{x}_{t+\lambda\times w}$.
Clearly, Window Switch changes only the order of list entries scanned and does not affect the total number of computations. Therefore, the time complexity of distance computation remains $\mathcal{O}\!\left(\frac{\|q\|}{s}\right)$.
@@
2 $\mathcal{X}\leftarrow\{\vec{x}_{i}\mid x_{i}^{j}\in x_{i},\vec{x}_{i}\in\mathcal{D}\}$;
@@ 3.3.2. Construction and Search
The construction process of the full-precision Sindi index with Window Switch is detailed in Algorithm 1. Given a sparse vector dataset $\mathcal{D}$ with maximum dimension $d$. For each dimension $j$ (Line 1), all vectors $\vec{x}_{i}$ in $\mathcal{D}$ with $x_{i}^{j}$ in $x_{i}$ are collected into a temporary set $\mathcal{X}$(Line 2). These vectors are then appended into the corresponding windows of the inverted list $j$ based on their IDs (Lines 3-5). Finally, the constructed index $I$ is returned after processing all dimensions (Line 6). The time complexity of the construction process is $O(\|\mathcal{D}\|\|\bar{x}\|)$, where $\|\bar{x}\|$ represents the average number of non-zero entries, that is $\frac{\sum_{\vec{x}_{i}\in\mathcal{D}}{\|x_{i}\|}}{\|\mathcal{D}\|}$.
The search process for the full-precision Sindi index is summarized in Algorithm 2, consisting of three main stages: product computation, accumulation, and heap update. Given a query $\vec{q}$, an inverted index $I$, and the recall number of nearest neighbors $k$, the algorithm initializes a distance array $A$ with length equal to $\lambda$, setting all elements to zero (Lines 1–2), and creates an empty min-heap $H$ (Line 3). The outer loop iterates over all windows $w\in\{0,\dots,\sigma-1\}$ (Line 4). For each non-zero query component $q^{j}\in\vec{q}$ (Line 5), the algorithm performs SIMD-based batched multiplication between $q^{j}$ and all components $x_{i}^{j}$ contained in $I_{j,w}$, storing the results sequentially into the temporary product array $T^{j}$ (Line 6). Next, it retrieves each $x_{i}^{j}$ in $I_{j,w}$ (Lines 7-8), computes the mapped index $m$ of $A$ (Line 9), and accumulates $T^{j}[t]$ into $A[m]$ (Line 10). After processing all $q^{j}\in q$ for the current window, the algorithm proceeds to the heap update stage (Line 12). For each entry in $A$, if $A[m]$ is greater than the current minimum in $H$ or if the heap contains fewer than $k$ elements, the algorithm inserts the corresponding global vector ID $(m+\lambda\times w)$ along with its distance $A[m]$ into the heap (Lines 13–14). If the heap exceeds size $k$, the smallest element is removed (Lines 15–16). Finally, $A[m]$ is reset to zero to prepare for the next window (Line 17). After all windows have been processed, the heap $H$ contains at most $k$ vector IDs paired with their full-precision distances to the query, which is returned as the final result (Line 20).
Complexity. Assuming $l$ is the average of non-zero entries in the traversed list, that is $l=\frac{\sum_{q^{j}\in q}{I_{j}.size()}}{\|q\|}$. Even with Window Switch, the total number of non-zero entries traversed remains constant, which is $\|q\|\,l$. Therefore, the time complexity of a full-precision query is $\mathcal{O}(\frac{\|q\|\,l}{s})$. In conclusion, the computational cost of querying is independent of $\lambda$.
@@ 3.4. Analysis of Window Size’s Impact on Performance
While the Window Switch strategy does not change the overall computational complexity, it has a significant impact on memory access costs. When $\lambda$ decreases, the distance array becomes shorter, leading to a lower cache-miss rate during random writes. However, the number of windows $\sigma=\frac{\|\mathcal{D}\|}{\lambda}$ increases, causing more frequent random accesses when switching between inverted sub-lists. Therefore, an appropriate choice of $\lambda$ is required to balance these two effects. The following example can illustrate this:
@@
Figure 5 reports experimental results for the full-precision Sindi on the SPLADE-1M and SPLADE-FULL datasets. For each dataset, we executed 6,9806,\!980 queries under different window sizes $\lambda$ and measured the QPS. In addition, we used the Intel VTune Profiler to record memory bound metrics for two types of memory accesses: distance array updates and sub-list switches. Here, memory bound refers to the percentage of execution time stalled due to memory accesses. For the SPLADE-1M dataset, as $\lambda$ increases from 11K to 11M, the distance array miss rate decreases monotonically, while the sub-list switching miss rate increases monotonically. The total memory-bound latency reaches its minimum near $\lambda\approx 100$K, corresponding to the highest query throughput. The SPLADE-FULL dataset exhibits the same trend, confirming the existence of an optimal window size.
Based on this, the memory access latency for queries can be expressed in a double power-law form (bertsekas2021data, ; hennessy2011computer, ) as follows:
@@
  • The independent variable is the window size $\lambda$;
  • The term $A\lambda^{+\alpha}$ captures the increasing cost of distance array cache misses as $\lambda$ grows;
  • The term $B\lambda^{-\beta}$ reflects the decreasing cost of sub-list switching misses with larger $\lambda$,
  • The constant $C$ represents baseline memory access costs unrelated to $\lambda$.
According to the properties of the double power-law function, $T_{mem}(\lambda)$ reaches its minimum at $\lambda^{*}=\left(\frac{B\beta}{A\alpha}\right)^{\frac{1}{\alpha+\beta}}$. For small $\lambda\ll\lambda^{*}$, $T_{\text{mem}}$ is dominated by the sub-list switching term and decreases as $\lambda$ grows. For large $\lambda\gg\lambda^{*}$, the distance-array term dominates and $T_{\text{mem}}$ increases with $\lambda$. The optimum $\lambda=\lambda^{*}$ occurs when the two terms balance.
@@
Figure 5 shows that the dashed line represents the theoretical QPS curve. It’s derived by estimating $(\alpha,\beta)$ via log–log regression and $(A,B,C)$ via least-squares fitting of the double power-law model. For the SPLADE-1M dataset, the model predicts an optimal $\lambda^{*}\approx 7.35\times 10^{4}$, while for SPLADE-FULL it predicts $\lambda^{*}\approx 1.25\times 10^{5}$. These predictions are of the same order as the measured optimal $\lambda\approx 10^{5}$, with small deviations attributable to model abstraction and measurement variability. This agreement in both magnitude and trend supports the validity of our scaling analysis.
@@ 4. Approximate Inverted Index
We focus on optimizing the query process through pruning and re-ranking. Pruning reduces the size of vectors or lists to improve search efficiency, while re-ranking compensates for pruning-induced precision loss by computing full inner products for a select set of candidates. Together, these techniques achieve a significant performance boost with only a minimal sacrifice in accuracy.
@@ 4.1. Pruning Strategies
A notable advantage of sparse vectors is that a small number of high-valued non-zero entries can represent most of the information in the entire vector (splade, ). This property arises from the training mechanisms of sparse models. For example, Splade tends to concentrate important information in a few non-zero dimensions. From a semantic perspective, many low-valued non-zero entries correspond to stopwords (e.g., “is”, “the”), which can be pruned.
Let $\vec{x}_{i}^{\prime}$ denote the pruned version of document $\vec{x}_{i}$, and $\vec{q}^{\prime}$ the pruned version of query $\vec{q}$, $l$ is the average length of list before pruning, $l$’ is the average length of list after pruning. The reduction in computational cost achieved by pruning is
$\|q\|l-\|q^{\prime}\|l^{\prime}$
while the inner product error $\varepsilon$ introduced is
$\varepsilon=\sum_{\vec{x}_{i}\in\mathcal{D}}\Bigl{(}\delta(\vec{x}_{i},\vec{q})-\delta(\vec{x}^{\prime}_{i},\vec{q}^{\prime})\Bigr{)}.$
Smaller $\|x^{\prime}_{i}\|$ leads to higher throughput gains, but also larger inner product error. Therefore, pruning must be designed with a trade-off between efficiency and accuracy. The following example shows that it is possible to retain only part of the non-zero entries while incurring only a small loss in inner product accuracy.
@@
We prune each document vector $\vec{x}_{i}$ and the query vector $\vec{q}$ by retaining only the non-zero entries with the largest absolute values, producing $\vec{x}^{\prime}_{i}$ and $\vec{q}^{\prime}$. We vary the pruning ratio, defined as $\frac{\|x^{\prime}_{i}\|}{\|x_{i}\|}$, to control the proportion of entries preserved. Figure 6(a) shows the corresponding inner product error under different pruning ratios. The results indicate that the error decreases sharply as the ratio increases from $0.1$ to $0.3$, and becomes nearly zero when the ratio is between $0.5$ and $1$, revealing a saturation effect of pruning ratio on inner product error.
As discussed in Section 3, for full-precision Sindi the upper bound for computing the inner product between $\vec{x}_{i}$ and $\vec{q}$ is $\varTheta\!\left(\frac{\|q\|}{s}\right)$. For a given $\vec{q}$, the overall time complexity of the query process is $\mathcal{O}\!\left(\frac{\|q\|\,l}{s}\right)$, where $l$ is the average number of inverted lists traversed. Reducing query latency therefore amounts to reducing $l$, $\|x\|$, and $\|q\|$. This can be approached from three directions: pruning lists, pruning documents, and pruning queries. List pruning and document pruning are applied during the index-construction stage, while query pruning is applied at query time. In this work, we mainly focus on the construction stage, as query pruning and document pruning are essentially both forms of vector pruning. We compare three pruning strategies—list pruning, document-count pruning, and quality-ratio pruning—and analyze their respective advantages and disadvantages:
@@
List Pruning (LP). LP operates at the inverted-list level: for each dimension $j$, only the non-zero entries with the largest absolute values are retained in its inverted list $I_{j}$, restricting the list length to $l^{\prime}$. Since the size of $I_{j}$ varies across dimensions, highly valued $|x_{i}^{j}|$ entries in longer lists may be pruned, while lower-valued non-zero entries in shorter lists are retained. After this list-wise pruning, each document vector $\vec{x}_{i}$ is transformed into a pruned version $\phi_{\mathrm{LP}}(\vec{x}_{i})$ containing only those coordinates that survive the list truncation.
Vector Number Pruning (VNP). VNP reduces dimensionality at the vector level. Its core idea is to retain, for each vector $\vec{x}_{i}$, the $vn$ non-zero entries with the largest magnitudes, so that $\|x^{\prime}_{i}\|=vn$. However, since $\|x_{i}\|$ varies across vectors, this scheme cannot consistently preserve the non-zero entries that contribute the most to the inner product computation.
Vector Number Pruning (VNP). VNP applies the pruning operator $\phi_{\mathrm{VNP}}$ at the vector level. For each document vector $\vec{x}_{i}$, $\phi_{\mathrm{VNP}}(\vec{x}_{i})$ retains only the $vn$ non-zero entries with the largest absolute values, so that $\lVert\phi_{\mathrm{VNP}}(\vec{x}_{i})\rVert=vn$. Since $\lVert\vec{x}_{i}\rVert$ varies across vectors, this scheme cannot consistently preserve the non-zero entries that contribute the most to the inner product computation.
Mass Ratio Pruning (MRP). MRP applies the pruning operator $\phi_{\mathrm{MRP}}$ based on the cumulative sum of absolute values of a vector’s non-zero entries. For each document vector $\vec{x}_{i}$, $\phi_{\mathrm{MRP}}(\vec{x}_{i})$ ranks all non-zero entries in descending order of absolute value and retains the smallest prefix whose cumulative sum reaches a target fraction $\alpha$ of the vector’s total mass. This adaptive scheme discards low-value components that contribute little to the inner product while allowing vectors with different value-distribution to keep variable numbers of entries, reducing inverted list size without imposing a uniform length limit.
To formally introduce MRP, we first define the mass of a vector as the sum of the absolute values of its non-zero entries.
@@
$\xi(\vec{x})\;=\;\sum_{x^{j}\in x}\lvert x^{j}\rvert$.
Then we define the $\alpha$-mass subvector as the shortest prefix of sorted non-zero entries whose cumulative absolute value reaches a fraction $\alpha$ of the total mass.
@@
Consider a vector $\vec{x}\in\mathbb{R}^{d}$ and a permutation $\pi$ that orders the non-zero entries of $\vec{x}$ by non-increasing absolute value, i.e. $\lvert x^{\pi_{j}}\rvert\geq\lvert x^{\pi_{j+1}}\rvert$. For a constant $\alpha\in(0,1]$, let $1\leq r\leq\|x\|$ be the smallest integer satisfying
$\sum_{j=1}^{r}\lvert x^{\pi_{j}}\rvert\;\leq\;\alpha\,\xi(\vec{x}).$
The collection $x^{\prime}\_\alpha\;=\;\bigl{\{}x^{\pi_{j}}\bigr{\}}_{j=1}^{r}$ is the sparse representation of a vector $\vec{x}\_\alpha$, where $\vec{x}\_\alpha]$ is the $\alpha$-mass subvector of $\vec{x}$.
@@
Figure 7 illustrates three pruning methods applied to sparse vectors $\vec{x}_{1}$, $\vec{x}_{2}$, and $\vec{x}_{3}$. List Pruning prunes each list to size $l^{\prime}=2$, Vector Number Pruning retains $vn=2$ top entries of each vector, and Mass Ratio Pruning prunes each vector $\bar{x}_{i}$ to $\bar{x}_{i,\alpha}$ with $\alpha=0.7$. The pruning result is shown in the figure: (i) the three strategies all reduce the same computation, which is $\|q\|l-\|q^{\prime}\|l^{\prime}=9-6=3$; (ii) Mass Ratio Pruning’s inner product error is the smallest. The reason is that List Pruning fails to retain the larger value $x_{2}^{1}$, as each list is limited to a maximum of 2 vectors. Similarly, Vector Number Pruning does not retain $x_{3}^{3}$. In contrast, Mass Ratio Pruning minimizes error by prioritizing influential entries.
@@
4 $\vec{x}_{i}\_\alpha\leftarrow$ $\alpha$-mass vector of $\vec{x}_{i}$
5 $\mathcal{D^{\prime}}\leftarrow\mathcal{D^{\prime}}\,\cup\vec{x}_{i}\_\alpha$

### v2

@@ 3.3. Cache Optimization
When the dataset size $\mathcal{D}$ reaches the million scale, the distance array $A$ becomes correspondingly large. Random accesses to such a long array cause frequent cache misses, making query performance highly memory-bound. In addition, allocating a full-length distance array for every query incurs substantial memory overhead.
To address these issues, Sindi adopts the Window Switch strategy, which partitions the vector ID space into fixed-size windows and restricts each query to accessing IDs within a single window at a time. With a shorter distance array per window, the accessed entries are located within a much more compact memory region. This substantially improves spatial locality, and the resulting access pattern closely resembles a sequential scan, enabling hardware prefetching and reducing cache misses.
@@ 3.3.1. Window Switch
During index construction, Sindi partitions the dataset $\mathcal{D}$ into contiguous ID segments, referred to as windows. The window size is denoted by $\lambda$ ($0<\lambda\leq\lVert\mathcal{D}\rVert$), and the total number of windows is $\sigma=\left\lceil\frac{\lVert\mathcal{D}\rVert}{\lambda}\right\rceil$. The $w$-th window contains vectors from $\vec{x}_{w\lambda}$ to $\vec{x}_{(w+1)\lambda-1}$, and the window index of vector $\vec{x}_{i}$ is $\left\lfloor\tfrac{i}{\lambda}\right\rfloor$. Each inverted list $I_{j}$ is partitioned in the same way, so every list has $\sigma$ windows. We denote the $w$-th window of the $j$-th inverted list by $I_{j,w}$ ($0\leq w<\sigma$), whose entries are the non-zero $x_{i}^{j}$ in dimension $j$ for vectors in that ID range. Each $x_{i}^{j}$ implicitly carries the identifier $i$ through its subscript.
At query time, the length of the distance array $A$ is set to the window size $\lambda$, and all windows share this same $A$. Within a window, each vector $\vec{x}_{i}$ is mapped to a unique entry $A[i\bmod\lambda]$.
The search procedure for the $w$-th window proceeds in two steps:
(1) Inner product computation. For each scanned posting list $I_{j,w}$, compute the products $q^{j}\times x_{i}^{j}$ for all its entries, using SIMD instructions when possible. These products are written sequentially into a temporary array $T^{j}$, which has the same length as $I_{j,w}$ and is index-aligned with it — i.e., $T^{j}[t]$ stores the product for the $t$-th posting in $I_{j,w}$. The accumulation stage then adds each $T^{j}[t]$ into the corresponding $A[i\bmod\lambda]$ using $\mathcal{O}(1)$ time per update.
(2) Heap update. After processing all query dimensions for the current window, $A$ holds the final scores for that window’s candidates. Scan $A$ to insert the top-scoring vectors into a minimum heap $H$, which maintains the vector IDs and scores for the results to be returned. Here, each $A[t]$ corresponds to vector $\vec{x}_{t+\lambda\times w}$, so the global ID can be recovered from the local index $t$.
Note that Window Switch only changes the order of list entries scanned, without altering the number of arithmetic operations. Thus, the time complexity of distance computation remains $\mathcal{O}\!\left(\frac{\lVert q\rVert}{s}\right)$.
@@
2 $\mathcal{X}\leftarrow\{\,\vec{x}_{i}\in\mathcal{D}\mid x_{i}^{j}\neq 0\,\}$;
@@ 3.3.2. Construction and Search
The construction process of the full-precision Sindi index with Window Switch is shown in Algorithm 1. Given a sparse vector dataset $\mathcal{D}$ of dimension $d$, the algorithm iterates over each dimension $j$ (Line 1). For each $j$, it collects all vectors $\vec{x}_{i}\in\mathcal{D}$ having a non-zero entry $x_{i}^{j}$ into a temporary set $\mathcal{X}$ (Line 2). These vectors are then appended to the corresponding window $I_{j,w}$ of $I_{j}$ based on their IDs, where $w=\lfloor i/\lambda\rfloor$ (Lines 3–5). After processing all dimensions, the inverted index $I$ is returned (Line 6). The time complexity of the construction process is $\mathcal{O}(\lVert\mathcal{D}\rVert\lVert\bar{x}\rVert)$, where $\lVert\bar{x}\rVert=\frac{\sum_{\vec{x}_{i}\in\mathcal{D}}\lVert x_{i}\rVert}{\lVert\mathcal{D}\rVert}$ denotes the average number of nonzero entries per vector.
The search process for the full-precision Sindi index is summarized in Algorithm 2, and consists of three stages: product computation, accumulation, and heap update. Given a query $\vec{q}$, inverted index $I$, and target recall size $k$, a distance array $A$ of length $\lambda$ is initialized to zeros (Lines 1–2) and an empty min-heap $H$ is created (Line 3). The outer loop traverses all windows $w\in\{0,\dots,\sigma-1\}$ (Line 4). For each non-zero query component $q^{j}$ (Line 5), SIMD-based batched multiplication is performed with all $x_{i}^{j}$ in $I_{j,w}$, and the results are stored sequentially into a temporary product array $T^{j}$ aligned with $I_{j,w}$ (Line 6). Each $x_{i}^{j}$ is then retrieved (Lines 7–8), its mapped index $m=i\bmod\lambda$ computed (Line 9), and $T^{j}[t]$ accumulated into $A[m]$ (Line 10). After all $q^{j}$ for the current window are processed, the heap update stage begins (Line 12): each $A[m]$ that exceeds the heap minimum or when $H$ has fewer than $k$ entries is inserted into $H$ with its global ID $(m+\lambda w)$ and score $A[m]$ (Lines 13–14), removing the smallest if size exceeds $k$ (Lines 15–16). $A[m]$ is then reset to zero for the next window (Line 17). When all windows are processed, $H$ contains up to $k$ vector IDs with their full-precision distances to $\vec{q}$, which are returned as the final result (Line 20).
Complexity. Let $l=\frac{\sum_{q^{j}\in q}|I_{j}|}{\lVert q\rVert}$ denote the average number of nonzero entries in the traversed lists. With Window Switch, the total number of postings visited is still $\lVert q\rVert\,l$, so the time complexity of a full-precision query is $\mathcal{O}\!\left(\frac{\lVert q\rVert\,l}{s}\right)$, where $s$ is the SIMD batch size. Hence, the computational cost is independent of the window size $\lambda$.
@@ 3.4. Analysis of Window Size’s Impact on Performance
While the Window Switch strategy does not alter the overall computational complexity, two types of memory-access costs are sensitive to the window size $\lambda$:
  • Random-access cost to the distance array. During accumulation, each partial score $T^{j}[t]$ is added to $A[i\bmod\lambda]$, producing a random write pattern over $A$. When $\lambda$ decreases, the length of $A$ becomes smaller, more of it fits in cache, and this random-access cost decreases due to fewer cache misses.
  • Cache eviction to sub-list cost when switching windows. Under Window Switch, the search iterates over multiple posting sub-lists $I_{j,w}$ for different dimensions $j$ within the same window $w$. Switching from one dimension’s sub-list to another may evict previously cached sub-list data from memory. When $\lambda$ decreases, the number of windows $\sigma=\frac{\lVert\mathcal{D}\rVert}{\lambda}$ increases, leading to more frequent loading and eviction of these sub-lists, and hence increasing this cost.
Selecting $\lambda$ thus requires balancing the reduced random-access cost to the distance array against the increased cache-eviction cost to inverted sub-lists. The following example illustrates this trade-off:
@@
Figure 5 reports experimental results for the full-precision Sindi on the SPLADE-1M and SPLADE-FULL datasets. For each dataset, we executed 6,9806,\!980 queries under different window sizes $\lambda$ and measured the QPS. We also used the Intel VTune Profiler (Intel Corporation, 2025) to record memory bound metrics for two types of memory accesses: random accesses to the distance array (arising from accumulation writes) and cache eviction to sub-lists (occurring when switching dimensions within a window). Here, memory bound denotes the percentage of execution time stalled due to memory accesses. For the SPLADE-1M dataset, as $\lambda$ increases from 11K to 11M, the memory bound from distance array accesses increases monotonically, whereas that from sub-list cache evictions decreases monotonically. The total memory-bound latency reaches its minimum near $\lambda\approx 100$K, corresponding to the highest query throughput. The SPLADE-FULL dataset exhibits the same trend, confirming the existence of an optimal window size.
Based on this, the memory-access latency for queries can be modeled by a double power-law (Bertsekas and Gallager, 2021; Hennessy and Patterson, 2011):
@@
  • $\lambda$ is the window size;
  • $A\lambda^{+\alpha}$ models the increasing cost of random accesses to the distance array as $\lambda$ grows;
  • $B\lambda^{-\beta}$ models the decreasing cost of cache eviction to sub-lists with larger $\lambda$;
  • $C$ is the baseline memory-access cost unrelated to $\lambda$.
This function reaches its minimum at $\lambda^{*}=\left(\frac{B\beta}{A\alpha}\right)^{\frac{1}{\alpha+\beta}}$. When $\lambda\ll\lambda^{*}$, $T_{\text{mem}}$ is dominated by the sub-list eviction term and decreases as $\lambda$ increases; when $\lambda\gg\lambda^{*}$, the distance-array term dominates and $T_{\text{mem}}$ grows with $\lambda$. The optimum $\lambda^{*}$ occurs when these two terms are balanced.
@@
Figure 5 shows that the dashed line represents the theoretical QPS curve. It is obtained by estimating $(\alpha,\beta)$ via log–log regression (Draper, 1998) and $(A,B,C)$ via least-squares fitting (Lawson and Hanson, 1995) of the double power-law model. For the SPLADE-1M dataset, the model predicts an optimal $\lambda^{*}\approx 7.35\times 10^{4}$, while for SPLADE-FULL it predicts $\lambda^{*}\approx 1.25\times 10^{5}$. These predictions are reasonably consistent with the measured optimal $\lambda\approx 10^{5}$, with minor discrepancies attributable to hardware-level performance variability and measurement noise introduced by VTune profiling.
@@ 4. Approximate Inverted Index
This section focuses on optimizing the query process through pruning and reordering. In our framework, pruning serves as a form of coarse retrieval, reducing the number of non-zero entries or inverted-list lengths to quickly generate a compact candidate set. Reordering then plays the role of fine-grained ranking by computing exact inner products for this candidate set. Combining these two stages significantly improves query throughput while keeping the loss in recall negligible.
@@ 4.1. Pruning Strategies
A key property of sparse vectors is that a small number of high-valued non-zero entries can preserve the majority of the vector’s overall information content (Formal et al., 2021). This distribution pattern typically results from the training objectives of sparse representation models. For instance, SPLADE often concentrates the most informative components into a limited set of high-weight dimensions. Conversely, many low-valued non-zero entries correspond to common stopwords (e.g., “is”, “the”), which can be safely removed with minimal impact on retrieval quality.
Let $\vec{x}_{i}^{\prime}$ denote the pruned version of document $\vec{x}_{i}$, and $\vec{q}^{\prime}$ the pruned version of query $\vec{q}$. Let $l$ be the average posting-list length before pruning and $l^{\prime}$ the average posting–list length after pruning. The reduction in computational cost achieved by pruning is
$\|q\|\cdot l\;-\;\|q^{\prime}\|\cdot l^{\prime},$
For a given pruning operator $\phi$, we define the inner product error for document $\vec{x}_{i}$ as
$e^{(\phi)}_{i}=\delta(\vec{x}_{i},\vec{q})-\delta\!\bigl{(}\phi(\vec{x}_{i}),\phi(\vec{q})\bigr{)},$
where $\delta(\cdot,\cdot)$ denotes the exact inner product and $\phi(\cdot)$ applies the pruning transformation. The total inner product error over the dataset $\mathcal{D}$ is then
$\varepsilon^{(\phi)}=\sum_{\vec{x}_{i}\in\mathcal{D}}e^{(\phi)}_{i}$
Smaller $\|x^{\prime}_{i}\|$ yields higher query throughput, typically incurs a larger inner product error. Therefore, pruning must be designed with a trade-off between efficiency and accuracy. The following example shows that it is possible to retain a subset of high–value non-zero entries while incurring merely a small loss in inner product accuracy.
@@
Figure 6(a) reports the inner product error measured on a 100100K-scale dataset when retaining different proportions of the largest non-zero entries from both document and query vectors. The results show that the error drops rapidly as the retaining ratio increases from 0.1 to 0.3, and becomes almost negligible once the ratio exceeds 0.5. This indicates a saturation effect, where further increasing the retaining ratio yields minimal additional gains in accuracy.
As discussed in Section 3, for full-precision Sindi the upper bound for computing the inner product between $\vec{x}_{i}$ and $\vec{q}$ is $\varTheta\!\left(\frac{\|q\|}{s}\right)$. For a given $\vec{q}$, the overall query complexity is $\mathcal{O}\!\left(\frac{\|q\|\,l}{s}\right)$, where $l$ is the average number of inverted lists traversed. Thus, reducing query latency requires decreasing $l$, $\|x\|$, and $\|q\|$. This can be achieved via three approaches: list pruning, document pruning, and query pruning. List and document pruning are performed during index construction, while query pruning is applied at query time. In this work, we focus on the construction stage, since query pruning and document pruning are both forms of vector pruning. We compare three strategies—List Pruning (LP), Vector Number Pruning (VNP), and Mass Ratio Pruning (MRP)—and analyze their respective strengths and weaknesses.
@@
List Pruning (LP). LP operates at the inverted–list level: for each dimension $j$, it retains only the non–zero entries with the largest absolute values in $I_{j}$, limiting the list length to $l^{\prime}$. Since the size of $I_{j}$ varies across dimensions, some high–value $|x_{i}^{j}|$ entries in longer lists may be removed, while lower–value entries in shorter lists may be kept. After pruning, each document vector $\vec{x}_{i}$ becomes $\phi_{\mathrm{LP}}(\vec{x}_{i})$, containing only the coordinates that survive the list truncation.
Vector Number Pruning (VNP). VNP applies the pruning operator $\phi_{\mathrm{VNP}}$ at the vector level. For each document vector $\vec{x}_{i}$, $\phi_{\mathrm{VNP}}(\vec{x}_{i})$ retains the $vn$ non–zero entries with the largest absolute values, ensuring $\|\phi_{\mathrm{VNP}}(\vec{x}_{i})\|=vn$. Since $\|\vec{x}_{i}\|$ varies across vectors, high–value $|x_{i}^{j}|$ entries that contribute substantially to the inner product may still be removed under this fixed–size scheme.
Mass Ratio Pruning (MRP). MRP applies the pruning operator $\phi_{\mathrm{MRP}}$ based on the cumulative sum of the absolute values of a vector’s non-zero entries. For each document vector $\vec{x}_{i}$, $\phi_{\mathrm{MRP}}(\vec{x}_{i})$ ranks all non-zero entries in descending order of absolute value and retains the shortest prefix whose cumulative sum reaches a fraction $\alpha$ of the vector’s total mass. This adaptive scheme removes low–value entries that contribute little to the inner product, while allowing vectors with different value distributions to keep variable numbers of entries, thereby reducing inverted–list size without enforcing a uniform length limit.
To formally introduce MRP, we first define the mass of a vector.
@@
$mass(\vec{x})\;=\;\sum_{x^{j}\in x}\lvert x^{j}\rvert$.
We define the vector obtained after Mass Ratio Pruning as the $\alpha$-mass subvector.
@@
Let $\vec{x}\in\mathbb{R}^{d}$ and let $\pi$ be a permutation that orders the non-zero entries of $\vec{x}$ by non–increasing absolute value, i.e., $|x^{\pi_{j}}|\geq|x^{\pi_{j+1}}|$. For a constant $\alpha\in(0,1]$, let $1\leq r\leq\|x\|$ be the smallest integer satisfying
$\sum_{j=1}^{r}|x^{\pi_{j}}|\;\geq\;\alpha\times mass(\vec{x}).$
The $\alpha$-mass subvector, denoted $\alpha\text{-}mass(\vec{x})$, is the vector whose non-zero entries are $\{x^{\pi_{j}}\}_{j=1}^{r}$.
@@
Figure 7 illustrates the three pruning methods applied to sparse vectors $\vec{x}_{1}$, $\vec{x}_{2}$, and $\vec{x}_{3}$. List Pruning prunes each inverted list to size $l^{\prime}=2$, Vector Number Pruning retains the top $vn=2$ entries of each vector, and Mass Ratio Pruning prunes each $\vec{x}_{i}$ to its $\alpha\text{-}mass(\vec{x}_{i})$ with $\alpha=0.7$. The figure shows: (i) all three strategies result in the same reduction in computation, $\|q\|l-\|q^{\prime}\|l^{\prime}=9-6=3$; (ii) Mass Ratio Pruning yields the smallest inner product error. This is because List Pruning cannot retain the larger value $x_{2}^{1}$ when each list is limited to two vectors, and Vector Number Pruning drops $x_{3}^{3}$. In contrast, Mass Ratio Pruning prioritizes high–value entries, thereby minimizing error.
@@
4 $\vec{x}^{\prime}_{i}\leftarrow\alpha\text{-}mass(\vec{x}_{i})$
5 $\mathcal{D^{\prime}}\leftarrow\mathcal{D^{\prime}}\,\cup\vec{x}^{\prime}_{i}$